For a two-tailed test at the 0.01 significance level, what is the rejection region when a sample n=25 is drawn from a normal population with unknown standard deviation?

Select one: a. Less than -2.797 and greater than 2.797. b. Less than -2.492 and greater than +2.492. c. Less than -2.797 or greater than +2.797. d. None of the suggested answers are correct. e. Less than -2.492 or greater than +2.492.
Answer #1

Based on the information provided, the significance level is , the degrees of freedom are n-1=25-1=24 and the critical value for a two-tailed test is .

The rejection region for this two-tailed test is

Hence the answer becomes:-

Let me know in the comments if anything is not clear. I will reply ASAP! Please do upvote if satisfied!